# Online Learning of Interaction Dynamics with Dual Model Predictive Control for Multi-Agent Systems Using Gaussian Processes

### v1

@@ Abstract
Decentralized control in complex and uncertain multi-agent scenarios requires careful consideration of the interactions between the agents. In this context, this paper proposes a dual model predictive control (MPC) method using Gaussian process (GP) models for multi-agent systems. While Gaussian process MPC (GP-MPC) has been shown to be effective in predicting the dynamics of other agents, current methods do not consider the influence of the control input on the covariance of the predictions, and, hence, lack the dual control effect. Therefore, we propose a dual MPC that directly optimizes the actions of the ego agent, and the belief of the other agents by jointly optimizing their state trajectories as well as the associated covariance while considering their interactions through a GP. We demonstrate our GP-MPC method in a simulation study on autonomous driving, showing improved prediction quality compared to a baseline stochastic MPC. Simulation results show that our GP-MPC can learn the interactions between agents online, demonstrating the potential of GPs for dual MPC in uncertain and unseen scenarios.
@@ I Introduction
Many real-world multi-agent systems are complex and unsuitable for centralized control approaches and require distributed or decentralized controllers [1]. Examples include autonomous vehicles (AVs), robotics, aerial vehicles and energy systems, which often have coupled dynamics, constraints or policies, making these systems interactive. The control of a single agent requires the awareness and consideration of these interactions with other agents in its environment to realize the system’s safety and performance, without relying on communication. In the case of AVs, the control of such an agent includes maneuvering this agent and avoiding collisions with other agents. As AVs often lack explicit communication, we must deduce and consider these uncertain interactions when planning the AV’s motion [2].
Model predictive control (MPC) is an interesting paradigm for this local control as it can directly optimize the agent’s trajectories while considering safety constraints. Particularly, stochastic MPC (SMPC) can address uncertainties by incorporating a probabilistic description of model uncertainty into a stochastic optimal control problem (SOCP). In SMPC, the control inputs not only affect the system state, but also the probability distribution associated with this state, which is referred to as the dual control effect [3, 4]. This dual control effect is two-fold. Firstly, control policies with the dual control effect causally anticipate future observations, and, hence, will be less conservative in regions of low uncertainty. Secondly, dual control policies can feature an inherent probing effect that facilitates active learning [3]. Although solving the SOCP via the Bellman equation naturally yields these aspects of the dual control effect, this is often intractable and requires some approximation which may eliminate the dual control effect [4].
Learning-based MPC uses data-driven modeling and machine learning methods to improve the system model and the parameterization of the SOCP. In particular, Gaussian processes (GPs) are an often-used modeling approach for non-parametric learning-based MPC, as they can compensate for complex model mismatch and enable direct assessment of the approximate model uncertainty [5]. As GPs enable universal and differentiable regression [6], joint conditioning of states [7], and tractable uncertainty propagation [8], they are of great interest for learning-based and interaction-aware dual MPC.
@@ I-A Related Work
Dual MPC is shown to be effective in simultaneous identification and control of systems with parametric and structural uncertainties [9]. In addition, Dual Model Predictive Path Integral control, a sampling-based MPC, has recently demonstrated the potential of dual control and active learning in multi-agent environments [10]. However, these methods currently rely on parametric models with fixed model structures. Furthermore, GP-MPC has been used to learn and predict the motion of other agents in autonomous racing [11, 12], and autonomous driving [13]. However, [11] fixes these predictions prior to solving the MPC problem making the controller agnostic to the interactions. Moreover, [12, 11, 13] do not exploit the dual control effect, as they tighten the constraints based on the predicted covariance of the previous MPC solution, which is independent of the control input sequence to be optimized. While [12, 13] rely on offline training data, using online training data can improve the generalizability of learning-based control in unseen scenarios, as can be observed in [11].
@@ I-B Contributions
In this paper, we present a novel learning-based dual MPC method to learn the unknown dynamics and/or policy of other agents, and bridge the previously identified research gap through the following contributions: (i) We extend GP-MPC for the local control of a single agent in a multi-agent environment by including the dual control effect. The MPC leverages the joint probability of the GP to adapt the state covariance of the other agents through the dual control effect, using Bayesian inference to learn the interactions between the ego agent and other the agents. (ii) We apply the proposed method in a simulation study on lane merging with an autonomous vehicle. In this study, the GP-MPC learns the closed-loop policy of another agent. By collecting training data online, the GP-MPC can adapt its prediction model and handle uncertain as well as unseen behavior of other agents. To the authors’ best knowledge, this is the first work that uses GPs to learn and predict the interactions between agents using dual MPC.
@@ II Problem Formulation
$x^{0}_{k+1}=f^{0}(x^{0}_{k},u^{0}_{k})$ (1)
that is subject to state and input constraints, $x^{0}_{k}\in\mathbb{X}^{0}\subseteq\mathbb{R}^{n_{x}}$, $u^{0}_{k}\in\mathbb{U}^{0}\subseteq\mathbb{R}^{n_{u}}$ at discrete time $k\in\mathbb{N}$. We assume to have perfect knowledge of the function $f^{0}$. The dynamics of Agent $j\in\mathcal{J}$, where $\mathcal{J}=\{1,2,\dots,n_{a}-1\}$, are composed of a known, nominal function $f^{j}$, and, an unknown and uncertain, residual function $g^{j}$ that influences the state via the full column rank matrix $B^{j}\in\mathbb{R}^{n_{x}\times n_{g}}$:
$x^{j}_{k+1}=f^{j}\bigl{(}x^{j}_{k}\bigr{)}+B^{j}g^{j}\bigl{(}\mathbf{x}_{k},\mathbf{u}_{k}\bigr{)},$ (2)
where ${\mathbf{x}_{k}=\bigl{(}x_{k}^{0},x_{k}^{1},\dots,x_{k}^{n_{a}-1}\bigr{)}}$, and $\mathbf{u}_{k}=\bigl{(}u_{k}^{0},u_{k}^{1},\dots,u_{k}^{n_{a}-1}\bigr{)}$ are the joint state and input vector. Note that the residual dynamics $g^{j}$ depends on $\mathbf{x}_{k}$ and $\mathbf{u}_{k}$, making the agents interactive. We assume that the control input of Agent $j\in\mathcal{J}$ is composed of a closed-loop state feedback policy $\kappa^{j}(\mathbf{x}_{k})$ and a stationary zero-mean disturbance process $w^{j}_{k}$:
$u^{j}_{k}=\kappa^{j}(\mathbf{x}_{k})+w^{j}_{k}.$ (3)
As Agent 0 shares its environment with Agent $j\in\mathcal{J}$, we have to consider their coupled safety constraints that define a safe set ${\mathbb{X}^{j}_{k}\bigl{(}x^{j}_{k}\bigr{)}=\{x^{0}_{k}\in\mathbb{R}^{n_{x}}\,\mid\,h^{j}\bigl{(}x^{0}_{k},x^{j}_{k}\bigr{)}\leq 0\}}$. We assume that $f$, $g$ and $h$ are continuously differentiable functions. We consider the problem of finding a local controller that safely controls Agent 00 in the sense that $x^{0}_{k}\in\bigcap_{j\in\mathcal{J}}\mathbb{X}^{j}_{k}\bigl{(}x^{j}_{k}\bigr{)}\cap\mathbb{X}^{0}$.
@@ III-A Modeling Interactions Between Agents
In this section, we propose a learning-based dual MPC for the control of Agent 00. Since we do not have full knowledge of the dynamics of Agent $j\in\mathcal{J}$, nor of its policy, it is essential that we consider the uncertainty of the residual dynamics $g^{j}$ in our MPC formulation in order to realize the safety of the controller as these interactions will affect the safe set $\mathbb{X}^{j}_{k}(x^{j}_{k})$.
As we only have partial information of the other agents, we predict their future states through Bayesian inference using GP regression. Firstly, we collect measurements of the states of all agents and the ego’s input $\mathbf{z}_{k}=[\mathbf{x}_{k}^{\top},u_{k}^{0\top}]^{\top}\in\mathbb{R}^{n_{z}}$, and we collect samples of the residual dynamics $y^{j}_{k}\in\mathbb{R}^{n_{g}}$ as follows:
$y^{j}_{k}=g^{j}\bigl{(}\mathbf{x}_{k},\mathbf{u}_{k}\bigr{)}=B^{j^{\dagger}}\bigl{(}x_{k+1}^{j}-f^{j}(x_{k})\bigr{)},$ (4)
where $B^{j^{\dagger}}=\bigl{(}B^{j\top}B^{j}\bigr{)}^{-1}B^{j\top}$. Subsequently, we use the collected training data $\mathbf{Z}=[\mathbf{z}_{1},\mathbf{z}_{2},\dots,\mathbf{z}_{n_{D}}]\in\mathbb{R}^{n_{z}\times n_{D}}$ and ${\mathbf{y}^{j}=[y_{1}^{j},y_{2}^{j},\dots,y_{n_{D}}^{j}]\in\mathbb{R}^{n_{g}\times n_{D}}}$ for inference by approximating the residual dynamics by a GP $\mathbf{d}^{j}:\mathbb{R}^{n_{z}}\rightarrow\mathbb{R}^{n_{g}}$ such that:
$g^{j}\bigl{(}\mathbf{x}_{k},\mathbf{u}_{k}\bigr{)}\approx\mathbf{d}^{j}\left(\mathbf{z}_{k}\right).$ (5)
In order to predict the future states of the other agents, we condition the posterior of the GP on the training data $\mathcal{D}=\bigl{\{}\mathbf{Z},\bigl{(}\mathbf{y}^{j}\bigr{)}_{j=1}^{n_{a}-1}\bigr{\}}$. For the sake of simplicity, we confine ourselves to scalar GPs $\left(n_{g}=1\right)$. Typically, multiple variables are predicted with $n_{g}\leq n_{x}$ independent GPs. We refer the interested reader to [8], for details on vectorial GPs. As we focus on learning the interactions between agents, we assume that $\mathbf{Z}$ and $\mathbf{y}^{j}$ are measured without being affected by measurement noise. Extending the GP regression to account for measurement noise is part of future work.
@@ III-B Gaussian Process Regression
Let us focus on predicting the dynamics of a single Agent $j\ \in\mathcal{J}$ and omit the superscript $j$ for the sake of readability. Firstly, we impose a prior distribution on the GP $d$ through a user-defined kernel function $k\left(\mathbf{z},\mathbf{z}^{\prime}\right)$ [6]. As we use the GP for model augmentation, we employ a zero mean prior on the GP. We assume the measured training outputs $\mathbf{y}$, and the prior distribution of the GP $d\left(\mathbf{z}\right)$ to be jointly Gaussian:
$\begin{bmatrix}\mathbf{y}\\ d\left(\mathbf{z}\right)\end{bmatrix}\sim\mathcal{N}\left(\mathbf{0},\begin{bmatrix}K_{\mathbf{ZZ}}&\mathbf{k}_{\mathbf{Zz}}\\ \mathbf{k}_{\mathbf{zZ}}&k_{\mathbf{zz}}\end{bmatrix}\right),$ (6)
where $k_{\mathbf{zz}^{\prime}}=k\left(\mathbf{z},\mathbf{z}^{\prime}\right)$, $\mathbf{k}_{\mathbf{Zz}}\in\mathbb{R}^{{n_{D}}}$ is the concatenation of the kernel function evaluated at the test point $\mathbf{z}$ and the training set $\mathbf{Z}$, where $[\mathbf{k}_{\mathbf{Zz}}]_{i}=k\left(\mathbf{z}_{i},\mathbf{z}\right)$ and ${\mathbf{k}_{\mathbf{Zz}}^{\top}=\mathbf{k}_{\mathbf{zZ}}}$, and $K_{\mathbf{ZZ}}\in\mathbb{R}^{{n_{D}}\times{n_{D}}}$ is a Gramm matrix, which satisfies ${[K_{\mathbf{ZZ^{\prime}}}]_{ij}=k\left(\mathbf{z}_{i},\mathbf{z}^{\prime}_{j}\right)}$. Marginalizing the jointly Gaussian prior over the training outputs $\mathbf{y}$ yields the GP posterior [6]:
$\text{Pr}\bigl{(}d\mid\mathbf{z},\mathcal{D}\bigr{)}=\mathcal{N}\bigl{(}\mu^{d}\left(\mathbf{z}\right),\Sigma^{d}\left(\mathbf{z}\right)\bigr{)},$ (7)
@@
$\mu^{d}(\mathbf{z})$ $=\mathbf{k}_{\mathbf{zZ}}K^{-1}_{\mathbf{Z}}\mathbf{y},$ (8a)
$\Sigma^{d}(\mathbf{z})$ $=k_{\mathbf{zz}}-\mathbf{k}_{\mathbf{zZ}}K^{-1}_{\mathbf{Z}\mathbf{Z}}\mathbf{k}_{\mathbf{Z}\mathbf{z}}.$ (8b)
Subsequently, we query the posterior of the GP $d^{j}(\mathbf{z})$, at a test point $\mathbf{z}\in\mathbb{R}^{n_{z}}$, to predict the residual dynamics $g^{j}$ over a time prediction horizon $i=0,1,\dots,N-1$:
@@
where $x_{i\mid k}$ denotes the $i$-th future prediction of $x$ made at time $k$, and $\mu^{d^{j}}$ and $\Sigma^{d^{j}}$ denote the mean and covariance function of the posterior distribution of the GP at a given test point $\mathbf{z}$.

### v2

@@ Abstract
The control of a single agent in complex and uncertain multi-agent environments requires careful consideration of the interactions between the agents. In this context, this paper proposes a dual model predictive control (MPC) method using Gaussian process (GP) models for multi-agent systems. While Gaussian process MPC (GP-MPC) has been shown to be effective in predicting the dynamics of other agents, current methods do not consider the influence of the control input on the covariance of the predictions, and hence lack the dual control effect. Therefore, we propose a dual MPC that directly optimizes the actions of the ego agent, and the belief of the other agents by jointly optimizing their state trajectories as well as the associated covariance while considering their interactions through a GP. We demonstrate our GP-MPC method in a simulation study on autonomous driving, showing improved prediction quality compared to a baseline stochastic MPC. The results show that GP-MPC can learn the interactions between the agents online, demonstrating the potential of GPs for dual MPC in uncertain and unseen scenarios.
@@ I Introduction
Many real-world multi-agent systems are complex and unsuitable for centralized control approaches and require distributed or decentralized controllers [1]. Examples include autonomous vehicles (AVs), robotics, aerial vehicles and energy systems, which often have coupled dynamics, constraints or policies, making these systems interactive. The control of a single agent requires the awareness and consideration of these interactions with other agents in its environment to achieve safety and performance without relying on communication between the agents. In the case of AVs, the control of such an agent includes maneuvering and avoiding collisions with other agents. As AVs often lack explicit communication, we must deduce and consider these uncertain interactions when planning the AV’s motion [2].
Model predictive control (MPC) is an interesting paradigm for this local control as it can directly optimize the agent’s trajectories while ensuring safety constraints are satisfied. Particularly, stochastic MPC (SMPC) can address uncertainties by incorporating a probabilistic description of model uncertainty into a stochastic optimal control problem (SOCP). In SMPC, the control inputs not only affect the system state, but also the probability distribution associated with this state, which is referred to as the dual control effect [3, 4]. This dual control effect is two-fold. Firstly, control policies with the dual control effect causally anticipate future observations, and, hence, will be less conservative in regions of low uncertainty. Secondly, dual control policies can feature an inherent probing effect that facilitates active learning [3]. Although solving the SOCP via the Bellman equation naturally yields these aspects of the dual control effect, this is often intractable and requires some approximation which may eliminate the dual control effect [4]. In this work, we propose a learning-based MPC method aimed towards advancing dual MPC in complex and dynamic environments.
Learning-based MPC uses data-driven modeling and machine learning methods to improve the system model and the parameterization of the SOCP. In particular, Gaussian processes (GPs) are an often-used modeling approach for non-parametric learning-based MPC, as they can compensate for complex model mismatch and enable direct assessment of the approximate model uncertainty [5]. As GPs enable universal and differentiable regression [6], joint conditioning of states [7], and tractable uncertainty propagation [8], they have strong potential for interaction-aware dual MPC.
@@ I-A Related Work
Dual MPC is shown to be effective in simultaneous identification and control of systems with parametric and structural uncertainties [9]. In addition, sampling-based methods have recently demonstrated the potential of interaction-aware control [10], and active learning in multi-agent environments [11]. However, these methods currently rely on parametric models with fixed model structures. By extending these methods to non-parametric models, a larger class of problems can potentially be addressed. Gaussian process MPC has been used to learn and predict the motion of other agents in autonomous racing [12, 13], and autonomous driving [14]. However, previous works fix the predictions of the other agents [12] or the covariance of these predictions [13, 12, 14] prior to solving the MPC problem, making the controller agnostic to the interactions. Consequently, these methods lack the dual control effect. While [13, 14] rely on offline training data, using online training data can help to reduce uncertainty and to improve the generalizability of learning-based control in unseen scenarios, as can be seen in [12].
@@ I-B Contributions
In this paper, we present a novel learning-based dual MPC method to learn the unknown dynamics and/or policy of other agents and bridge the previously identified research gap through the following contributions: (i) We extend GP-MPC for the local control of a single agent in a multi-agent environment by including the dual control effect through a custom GP implementation. Although we do not incentivize active learning, our MPC explicitly considers the joint covariance of the agents in its optimization. The MPC leverages the joint probability of the GP to adapt the state covariance of the other agents through the dual control effect, using Bayesian inference to predict the interactions between the ego agent and the other agents. (ii) Our GP-MPC uses online measurements to update the GP. Through online learning, the GP-MPC can adapt its predictions and handle uncertain and unseen behavior of other agents. (iii) We apply the proposed method in a simulation study on lane merging with an autonomous vehicle. In this study, the GP-MPC learns the closed-loop policy of another agent and adapts the constraints based on its confidence. To the authors’ best knowledge, this is the first work that uses GPs to learn and predict the interactions between agents online via dual MPC.
@@ II Problem Formulation
$x^{0}_{k+1}=f^{0}(x^{0}_{k},u^{0}_{k})\vspace{-2pt}$ (1)
that is subject to state and input constraints, $x^{0}_{k}\in\mathbb{X}^{0}\subseteq\mathbb{R}^{n_{x}}$, $u^{0}_{k}\in\mathbb{U}^{0}\subseteq\mathbb{R}^{n_{u}}$ at discrete time step $k\in\mathbb{N}$. We assume to have perfect knowledge of the function $f^{0}$. The dynamics of Agent $j\in\mathcal{J}$, where $\mathcal{J}=\{1,2,\dots,n_{a}-1\}$, are composed of a known, nominal function $f^{j}$, and, an unknown and uncertain, residual function $g^{j}$ that influences the state via the full column rank matrix $B^{j}\in\mathbb{R}^{n_{x}\times n_{g}}$:
$x^{j}_{k+1}=f^{j}\bigl{(}x^{j}_{k}\bigr{)}+B^{j}g^{j}\bigl{(}\mathbf{x}_{k},\mathbf{u}_{k}\bigr{)}+w^{j}_{k},\vspace{-2pt}$ (2)
where ${\mathbf{x}_{k}=\mathrm{vec}\bigl{(}x_{k}^{0},x_{k}^{1},\dots,x_{k}^{n_{a}-1}\bigr{)}}$ is the joint state vector, and $\mathbf{u}_{k}=\mathrm{vec}\bigl{(}u_{k}^{0},u_{k}^{1},\dots,u_{k}^{n_{a}-1}\bigr{)}$ is the joint input vector, and $w^{j}_{k}\in\mathbb{R}^{n_{x}}$ is a Gaussian disturbance process. Note that the residual dynamics $g^{j}$ depends on $\mathbf{x}_{k}$ and $\mathbf{u}_{k}$, making the agents interactive. We assume that the control input of Agent $j\in\mathcal{J}$ is composed of a closed-loop state feedback policy:
$u^{j}_{k}=\kappa^{j}(\mathbf{x}_{k}).\vspace{-2pt}$ (3)
As Agent 0 shares its environment with Agent $j\in\mathcal{J}$, we have to consider their coupled safety constraints that define a safe set ${\mathbb{X}^{j}_{k}\bigl{(}x^{j}_{k}\bigr{)}=\{x^{0}_{k}\in\mathbb{R}^{n_{x}}\,\mid\,h^{j}\bigl{(}x^{0}_{k},x^{j}_{k}\bigr{)}\leq 0\}}$. We assume that $f$, $g$, $h$ and $\kappa$ are continuously differentiable functions. We consider the problem of finding a controller that safely controls Agent 00 in the sense that $x^{0}_{k}\in\bigcap_{j\in\mathcal{J}}\mathbb{X}^{j}_{k}\bigl{(}x^{j}_{k}\bigr{)}\cap\mathbb{X}^{0}$.
@@ III-A Modeling Interactions Between Agents
In this section, we propose a learning-based dual MPC for the control of Agent 00. Since we do not have full knowledge of the dynamics of Agent $j\in\mathcal{J}$, nor of its policy, i.e., of its intentions, it is essential that we consider the uncertainty of the residual dynamics $g^{j}$ and their policy $\kappa^{j}$ in the formulation of our MPC formulation in order to realize the safety of the controller as these interactions will affect the safe set $\mathbb{X}^{j}_{k}(x^{j}_{k})$. As we only have partial information of the other agents, we predict their future states through Bayesian inference using GP regression. Firstly, we collect measurements of the states of all agents and the ego’s input $\mathbf{z}_{k}=[\mathbf{x}_{k}^{\top},u_{k}^{0\top}]^{\top}\in\mathbb{R}^{n_{z}}$, and we collect samples of the residual dynamics $y^{j}_{k}\in\mathbb{R}^{n_{g}}$ as follows:
$y^{j}_{k}=B^{j^{\dagger}}\bigl{(}x_{k+1}^{j}-f^{j}(x_{k})\bigr{)}=g^{j}\bigl{(}\mathbf{x}_{k},\kappa^{j}(\mathbf{x}_{k}),u^{0}_{k}\bigr{)}+\tilde{w}^{j}_{k},\vspace{-1pt}$ (4)
with $B^{j^{\dagger}}=\bigl{(}B^{j\top}B^{j}\bigr{)}^{-1}B^{j\top}$ and $\tilde{w}^{j}_{k}\sim\mathcal{N}(0,\Sigma^{j}_{w})$. Then, we use the collected training data $\mathbf{Z}=[\mathbf{z}_{1},\mathbf{z}_{2},\dots,\mathbf{z}_{n_{D}}]\in\mathbb{R}^{n_{z}\times n_{D}}$, ${\mathbf{y}^{j}=[y_{1}^{j},y_{2}^{j},\dots,y_{n_{D}}^{j}]\in\mathbb{R}^{n_{g}\times n_{D}}}$ in order to approximate the residual dynamics by a GP $\mathbf{d}^{j}:\mathbb{R}^{n_{z}}\rightarrow\mathbb{R}^{n_{g}}$ such that:
$g^{j}\bigl{(}\mathbf{x}_{k},\kappa^{j}(\mathbf{x}_{k}),u^{0}_{k}\bigr{)}\approx\mathbf{d}^{j}\left(\mathbf{z}_{k}\right).\vspace{-2pt}$ (5)
In order to predict the future states of the other agents, we condition the posterior of the GP on the training data $\mathcal{D}=\bigl{\{}\mathbf{Z},\bigl{(}\mathbf{y}^{j}\bigr{)}_{j=1}^{n_{a}-1}\bigr{\}}$. For the sake of simplicity, we confine ourselves to scalar GPs $\left(n_{g}=1\right)$. Typically, multiple variables are predicted with $n_{g}\leq n_{x}$ independent GPs. We refer the interested reader to [8], for details on vectorial GPs.
@@ III-B Gaussian Process Regression
Let us focus on predicting the dynamics of a single Agent $j\ \in\mathcal{J}$ and omit the superscript $j$ for the sake of readability. Firstly, we impose a prior distribution on the GP $d\left(\mathbf{z}\right)\in\mathbb{R}$ through a user-defined kernel function $k\left(\mathbf{z},\mathbf{z}^{\prime}\right)$ [6]. As we use the GP for model augmentation, we employ a zero mean prior. We assume that the measured training outputs $\mathbf{y}$, and the prior distribution of the GP $d\left(\mathbf{z}\right)$ are jointly Gaussian:
$\begin{bmatrix}\mathbf{y}\\ d\left(\mathbf{z}\right)\end{bmatrix}\sim\mathcal{N}\left(\mathbf{0},\begin{bmatrix}K_{\mathbf{ZZ}}+I\sigma_{w}^{2}&\mathbf{k}_{\mathbf{Zz}}\\ \mathbf{k}_{\mathbf{zZ}}&k_{\mathbf{zz}}s^{2}\end{bmatrix}\right),$ (6)
where $k_{\mathbf{zz}^{\prime}}=k\left(\mathbf{z},\mathbf{z}^{\prime}\right)$, $\mathbf{k}_{\mathbf{Zz}}\in\mathbb{R}^{{n_{D}}}$ is the concatenation of the kernel function evaluated at the test point $\mathbf{z}$ and the training set $\mathbf{Z}$, where $[\mathbf{k}_{\mathbf{Zz}}]_{i}=k\left(\mathbf{z}_{i},\mathbf{z}\right)$ and ${\mathbf{k}_{\mathbf{Zz}}^{\top}=\mathbf{k}_{\mathbf{zZ}}}$, and $K_{\mathbf{ZZ}}\in\mathbb{R}^{{n_{D}}\times{n_{D}}}$ is a Gram matrix, which satisfies ${[K_{\mathbf{ZZ^{\prime}}}]_{ij}=k\left(\mathbf{z}_{i},\mathbf{z}^{\prime}_{j}\right)}$. The Gaussian noise on $\mathbf{y}\in\mathbb{R}^{1\times n_{D}}$ is typically handled by regularizing the Gram matrix with the noise variance $\sigma_{w}^{2}$, while the predictions on $d$ are noise free. Marginalizing the jointly Gaussian prior over the training outputs $\mathbf{y}$ yields the GP posterior [6], which is our conditioned belief of $d(\mathbf{z})$ due to the observations $\mathcal{D}$:
$\text{Pr}\bigl{(}d(\mathbf{z})\mid\mathbf{z},\mathcal{D}\bigr{)}=\mathcal{N}\bigl{(}\mu^{d}\left(\mathbf{z}\right),\Sigma^{d}\left(\mathbf{z}\right)\bigr{)},\vspace{-2pt}$ (7)
@@
$\mu^{d}(\mathbf{z})$ $=\mathbf{k}_{\mathbf{zZ}}\left(K_{\mathbf{ZZ}}+I\sigma^{2}_{w}\right)^{-1}\mathbf{y},$ (8a)
$\Sigma^{d}(\mathbf{z})$ $=k_{\mathbf{zz}}-\mathbf{k}_{\mathbf{zZ}}\left(K_{\mathbf{ZZ}}+I\sigma^{2}_{w}\right)^{-1}\mathbf{k}_{\mathbf{Z}\mathbf{z}}.$ (8b)
Subsequently, we query the posterior of the GP $d^{j}(\mathbf{z})$, at a test point $\mathbf{z}\in\mathbb{R}^{n_{z}}$, to predict the residual dynamics $g^{j}$ of Agent $j$ over a time prediction horizon $i=0,1,\dots,N-1$:
@@
where $x_{i\mid k}$ denotes the $i$-th step-ahead prediction of $x_{k+i}$ made at time step $k$ given the data $\mathcal{D}_{k}$, where $\mu^{d^{j}}$ and $\Sigma^{d^{j}}$ denote the predicted mean and covariance function of the posterior distribution of the GP at a given test point $\mathbf{z}$.